# Break-The-Chain: Reasoning Failures in LLMs via Adversarial Prompting in Code Generation

### v1

@@ Abstract
Large Language Models (LLMs) have achieved remarkable success in tasks requiring complex reasoning, such as code generation, mathematical problem solving, and algorithmic synthesis — especially when aided by reasoning tokens and Chain-of-Thought prompting. Yet, a core question remains: do these models truly reason, or do they merely exploit shallow statistical patterns? In this paper, we systematically investigate the robustness of reasoning LLMs by introducing a suite of semantically faithful yet adversarially structured prompt perturbations. Our evaluation — spanning 700 perturbed code generations derived from LeetCode-style problems — applies transformations such as storytelling reframing, irrelevant constraint injection, example reordering, and numeric perturbation. We observe that while certain modifications severely degrade performance (with accuracy drops up to -42.1%), others surprisingly improve model accuracy by up to 35.3%, suggesting sensitivity not only to semantics but also to surface-level prompt dynamics. These findings expose the fragility and unpredictability of current reasoning systems, underscoring the need for more principles approaches to reasoning alignments and prompting robustness. We release our perturbation datasets and evaluation framework  [^1] https://github.com/jrohsc/Break-the-Chain to promote further research in trustworthy and resilient LLM reasoning.
Figure 1: Overview of Break-the-Chain
@@ 1 Introduction (Problem Statement)
In this work, we extend this line of inquiry to the domain of code generation. We introduce BREAK-THE-CHAIN, a systematic framework for evaluating reasoning robustness in LLMs using semantically aligned yet adversarial prompt perturbations. Our perturbations span a diverse set of transformations—including storytelling, gamification, domain shifts, distracting constraints, and negation—that vary in narrative structure, lexical framing, and semantic drift. Crucially, each transformation preserves the core logic of the original task, allowing us to isolate how surface-level changes impact reasoning fidelity.
@@ 2 Related work
The challenge of evaluating and ensuring true reasoning capabilities in Large Language Models (LLMs), particularly for complex tasks such as code generation, has been approached from several angles. While LLMs have demonstrated strong performance across multiple NLP tasks, significant challenges remain in ensuring their robustness, reliability, and the actual depth of their reasoning quality. Our work, BREAK-THE-CHAIN, builds upon and extends prior research in three key areas that inform our study on reasoning failures under adversarial prompt perturbations: the nature of reasoning in LLMs, specific advancements and evaluations in LLM-based code generation, and existing research on attacks against LLM reasoning.
@@ 2.1 Reasoning Large Language Models
Reasoning in LLMs encompasses analytical tasks like math problem-solving, code generation, and logical inference. While traditional models often rely on pattern recognition, reasoning-enhanced models increasingly use techniques such as reasoning tokens and Chain-of-Thought (CoT) prompting Wei et al., (2022). Reasoning tokens are designed as explicit markers within the training process to guide LLMs toward more structured reasoning, an approach that has shown improvements in handling complex reasoning challenges, especially in arithmetic and code-related tasks. CoT prompting Wei et al., (2022) and its derivatives, like Zero-shot-CoT Wang et al., (2023) and Program of Thoughts Chen et al., (2022) (which aims to separate computation from reasoning), have demonstrably enhanced multi-step reasoning. However, while CoT can improve performance on standard benchmarks, our work with BREAK-THE-CHAIN investigates a critical subsequent question: whether such induced reasoning chains remain stable and effective when the problem’s surface structure is perturbed, even if the core logic is preserved. This inquiry is crucial for understanding the true depth and transferability of the elicited reasoning.
Recent surveys and comprehensive reviews Mondorf and Plank, (2024); Prasad et al., (2023); Golovneva et al., (2022) have further explored the reasoning behavior of LLMs beyond simple accuracy metrics, highlighting a prevalent concern: LLMs often rely on surface-level correlations rather than genuine logical inference. For instance, models might associate certain keywords or phrasing patterns with specific solution templates. BREAK-THE-CHAIN directly tests this dependency by altering these patterns through methods like Storytelling, Gamification, and Domain Shift, all while maintaining the underlying semantic integrity of the task. This reliance on potentially superficial cues, even with advanced prompting techniques, underscores the need for evaluations like ours. In this work, we therefore aim to evaluate reasoning through more nuanced behavioral analyses rather than just final answer correctness, focusing on behavioral consistency under controlled linguistic and structural variations.
@@ 2.2 LLM for Code Generation
While these methods are valuable, metrics like pass@k primarily assess functional correctness on standard problem formulations, and LLM-based evaluators like CODEJUDGE, though assessing semantic correctness, may not fully capture how models respond when the presentation of the problem itself is adversarially yet semantically perturbed. Our study aligns with the goal of robust evaluation but extends it by focusing on the stability of the reasoning process leading to code generation when faced with such input variations. BREAK-THE-CHAIN specifically assesses the impact of these adversarial modifications on the reasoning tasks inherent in code generation, probing whether the logical process itself is resilient.
@@ 2.3 Reasoning LLM Attacks
Adversarial attacks on code generation models have also been investigated. The DeceptPrompt framework Wu et al., (2023) reveals that benign-looking natural language modifications can induce LLMs to generate insecure or incorrect code, exposing vulnerabilities in real-world applications. While DeceptPrompt reveals vulnerabilities leading to insecure code, BREAK-THE-CHAIN explores a complementary set of semantically-focused perturbations (e.g., Storytelling, Gamification, Domain Shift, Distracting Constraints). These are designed not primarily to induce insecurity, but to test the stability and fidelity of the underlying reasoning chain required for correct code generation when familiar problem cues are altered. Our study examines how these adversarial modifications—ranging from narrative restructuring to the injection of logically inert information—impact reasoning models. A key distinction in our methodology is the emphasis on perturbations that, while adversarial in their structure or framing, are designed to be semantically faithful to the original problem’s core logic. This allows us to isolate reasoning failures that are not due to misunderstanding a completely different task, but rather to an inability to reason robustly when the problem presentation changes.
In summary, prior work has established the potential of LLMs in complex reasoning and code generation, yet has also consistently highlighted their brittleness and potential over-reliance on learned superficial patterns. Existing evaluation methods and attack strategies provide valuable insights into model capabilities and vulnerabilities. However, a gap remains in systematically understanding how LLMs’ reasoning chains for code generation withstand diverse, semantically-grounded perturbations that alter problem presentation without changing the core logic. BREAK-THE-CHAIN addresses this by introducing a novel framework and a suite of perturbation techniques specifically designed to probe these nuanced aspects of reasoning robustness in the domain of code generation.
@@ 3.2 Perturbation Methods
This transformation wraps the task in a challenge-based format (e.g., player or robots solving the task) while keeping the logic intact. Similar to storytelling, it introduces anthropomorphic or thematic context, drawing inspiration from instruction-tuning setups that present tasks in game-like scenarios (Wu et al.,, 2023). We evaluate whether such gamified inputs boosts engagement or distracts models from formal reasoning (see Figure LABEL:fig_gamification_prompt).
@@
Inspired by the symbolic injection strategies in GSM-Symbolic (Mirzadeh et al.,, 2024), this method appends irrelevant but plausible-sounding constraints (e.g., ”input is a palindrome”) to the problem. While logically inert, these constraints test the model’s ability to filter out irrelevant linguistic noise. Unlike classic adversarial examples, these do not change task correctness but increase input ambiguity (see Figure LABEL:fig_distracting_constraints_prompt).
@@
This strategy replaces tokens and terminology from one context to another (e.g., ”array of integers: becomes ”daily revenue”) or replacing terms with domain-specific equivalents (e.g., convert numbers to elevator floors, jobs, floors, days) (see Figure LABEL:fig_domain_shift_prompt).
@@ 3.3 Logical Preservation Scoring
To quantify the degree of semantic deviation induced by each transformation, we developed a Logical Preservation Score. We employed the Claude-3.7-Sonnet model as an automated evaluator to ensure consistent and scalable scoring across all 700 perturbed instances. This approach mitigates potential human rater fatigue and subjectivity. The scoring is guided by a detailed rubric-based prompt (see Appendix LABEL:appendix:preservation, Figure LABEL:fig_logical_preservation_prompt) that instructs the model to rate preservation on a scale of 1-10, where 10 indicates perfect logical alignment and 1 indicates a complete semantic shift. These scores are used in later analysis to understand how semantic drift correlates with performance degradation.
@@ 5.1 Clean Performance Overview
We begin by benchmarking all models on the original (unmodified) version of 100 LeetCode-style problems (Table LABEL:tab_accuracy_clean). As expected, models optimized for instruction-following and reasoning, such as Gemini-2.5-Flash (95.0% overall) and Claude-3.7-Sonnet (90.0%), outperform code-specific models like Qwen2.5-Coder (59.0%) and DeepSeek-Coder-33B (54.0%). ON average, performance declines sharply from Easy (77.6%) to Hard (38.3%) tasks, reflecting the inherent challenge of generalization across reasoning complexity.
@@ 5.2 Perturbation Resilience and Gains
Table LABEL:tab_attack_table_transposed_sorted reveals that semantic perturbations do not uniformly degrade performance. On the contrary, several perturbation types — particularly Storytelling and Gamification — consistently improve performance, especially for models tuned for human-aligned reasoning.
@@ 5.3 Semantic Robustness
In Figure 3, we plot average model accuracy against our logical preservation scores (Table LABEL:tab_preservation_score. A positive, albeit nonlinear, correlation emerges: Storytelling (score: 8.89) and Gamification (8.01) lead to improved or stable performance. Distracting Constraints (3.82) and Negation Objective (1.83) consistently degrade it. However, we noticed that performance does not degrade proportionally with logical drift. Some logically preserved rewrites improve accuracy, while a single line of structurally trivial edits (like Distracting Constraints) dramatically hurt it. This suggests that cognitive alignment — not logical alignment alone — governs LLM success. LLMs are more sensitive to semantic distractors than to narrative abstraction.
@@ 5.5 Difficulty Analysis
As shown in Table LABEL:tab_reasoning_models_difficulty, models often retain high accuracy on Easy problems across all attacks, but break under Hard tasks even for high-preservation rewrites. This confirms that reasoning complexity, not just input phrasing, lead to model breakdown. However, storytelling notably boosts hard-problem performance for Gemini-2.5 (+11.7%) and LLaMA3.1 (+23.5%), indicating that human-friendly reformulations may facilitate deeper reasoning under challenge.
@@ 5.6 Errors through Prompt Clarity
This highlights a dual nature of perturbations: when they disrupt expected patterns without adding structure, they degrade performance; but when they elaborate or scaffold problem context, they may improve reasoning as we see above for certain models in Table LABEL:tab_code_models. Future models may benefit from tuning not only on correct answers, but on uncertainty aware reasoning paths.
@@ 5.7 Comment-Driven Reasoning Behaviors
A notable behavioral pattern emerged when analyzing Claude-3.7-Sonnet’s performance under the Gamification prompt variant. While Claude-3.7 demonstrates strong performance under clean conditions (90.0% overall accuracy; Table LABEL:tab_accuracy_clean), its accuracy drops dramatically by –40.0% under Gamification (Table LABEL:tab_attack_table_transposed_sorted). Manual inspection of model outputs (Table  LABEL:tab_claude_comment_example) revealed a consistent shift in behavior: Claude’s code responses on unperturbed prompts routinely included inline comments—clarifying control flow logic, edge case handling, or the purpose of specific variables. In contrast, responses under gamified prompt formulations lacked such explanatory commentary.
@@ 6.1 Methods
This method fully inverts the core task (e.g., ”maximize” $\rightarrow$ ”minimize” or ”include” $\rightarrow$ ”exclude”). A part of the transformation prompt asks the LLM to: ”rewrite it with the opposite objective … ensure the examples and output are updated accordingly: (see Figure LABEL:fig_negation_prompt). While the input, output, explanation and constraints are preserved, the solution logic must change, requiring the model to recompute from a new problem intent. The preservation score was recorded as 1.83 based on Claude-3.7.
@@
This variant applied milder semantic reversals (e.g., asking for ”non-decreasing” instead of ”increasing”) without flipping the task entirely. The instruction explicitly states: ”Do NOT fundamentally change the problem’s task … adjust the examples minimally” (see Figure LABEL:fig_soft_negation_prompt). This design tests whether models are robust to subtle linguistic tweaks that alter logic while retaining solvability. The preservation score was recorded as 2.56 based on Claude-3.7, a little higher than the hard negation.
@@ 6.2 Results
Table LABEL:tab_ablation summarizes the impact of negation-based perturbations. Hard negation proves highly disruptive, causing average accuracy drops of over 50% across all difficulty levels. For instance, Gemini-2.5-Flash drops from 95.0% to 42.6%, and Claude-3.7 falls from 90.0% to 15.0%. Even instruction-following models struggle to adapt, as reasoning reversal—not just language adaptation—is required. In contrast, soft negation shows more variability, with some models like LLaMA-3.1 improving from 19.0% to 29.4%. However, others like DeepSeek-14B and Claude-3.7 still suffer large drops, showing that even subtle logical shifts can disrupt reasoning.
@@ 8 Conclusion
In this work, we presented Break-The-Chain, a systematic framework for evaluating reasoning robustness in LLMs under adversarial prompting in code generation tasks. By introducing semantically plausible but logically challenging perturbations such as storytelling, gamification, distracting constraints, and negation objectives, we revealed key vulnerabilities in current state-of-the-art models. Through comprehensive experiments across nine LLMs and 700 perturbed instances, we found that while some perturbations surprisingly improved model performance, others, especially those involving subtle semantic drift, led to significant reasoning failures.

### v2

# Chain-of-Code Collapse: Reasoning Failures in LLMs via Adversarial Prompting in Code Generation

@@ Abstract
Large Language Models (LLMs) have achieved remarkable success in tasks requiring complex reasoning, such as code generation, mathematical problem solving, and algorithmic synthesis — especially when aided by reasoning tokens and Chain-of-Thought prompting. Yet, a core question remains: do these models truly reason, or do they merely exploit shallow statistical patterns? In this paper, we introduce Chain-of-Code Collapse, where we systematically investigate the robustness of reasoning LLMs by introducing a suite of semantically faithful yet adversarially structured prompt perturbations. Our evaluation — spanning 700 perturbed code generations derived from LeetCode-style problems — applies transformations such as storytelling reframing, irrelevant constraint injection, example reordering, and numeric perturbation. We observe that while certain modifications severely degrade performance (with accuracy drops up to -42.1%), others surprisingly improve model accuracy by up to 35.3%, suggesting sensitivity not only to semantics but also to surface-level prompt dynamics. These findings expose the fragility and unpredictability of current reasoning systems, underscoring the need for more principles approaches to reasoning alignments and prompting robustness. We release our perturbation datasets and evaluation framework  [^1] https://github.com/jrohsc/Chain-of-Code-Collapse to promote further research in trustworthy and resilient LLM reasoning.
Figure 1: Overview of Chain-of-Code Collapse
@@ 1 Introduction (Problem Statement)
In this work, we extend this line of inquiry to the domain of code generation. We introduce Chain-of-Code Collapse (CoCC), a systematic framework for evaluating reasoning robustness in LLMs using semantically aligned yet adversarial prompt perturbations. Our perturbations span a diverse set of transformations—including storytelling, gamification, domain shifts, distracting constraints, and negation—that vary in narrative structure, lexical framing, and semantic drift. Crucially, each transformation preserves the core logic of the original task, allowing us to isolate how surface-level changes impact reasoning fidelity.
@@ 2 Related work
The challenge of evaluating and ensuring true reasoning capabilities in Large Language Models (LLMs), particularly for complex tasks such as code generation, has been approached from several angles. While LLMs have demonstrated strong performance across multiple NLP tasks, significant challenges remain in ensuring their robustness, reliability, and the actual depth of their reasoning quality. Our work, builds upon and extends prior research in three key areas that inform our study on reasoning failures under adversarial prompt perturbations: the nature of reasoning in LLMs, specific advancements and evaluations in LLM-based code generation, and existing research on attacks against LLM reasoning.
@@ 2.1 Reasoning Large Language Models
Reasoning in LLMs encompasses analytical tasks like math problem-solving, code generation, and logical inference. While traditional models often rely on pattern recognition, reasoning-enhanced models increasingly use techniques such as reasoning tokens and Chain-of-Thought (CoT) prompting Wei et al., (2022). Reasoning tokens are designed as explicit markers within the training process to guide LLMs toward more structured reasoning, an approach that has shown improvements in handling complex reasoning challenges, especially in arithmetic and code-related tasks. CoT prompting Wei et al., (2022) and its derivatives, like Zero-shot-CoT Wang et al., (2023) and Program of Thoughts Chen et al., (2022) (which aims to separate computation from reasoning), have demonstrably enhanced multi-step reasoning. However, while CoT can improve performance on standard benchmarks, our work with CoCC investigates a critical subsequent question: whether such induced reasoning chains remain stable and effective when the problem’s surface structure is perturbed, even if the core logic is preserved. This inquiry is crucial for understanding the true depth and transferability of the elicited reasoning.
Recent surveys and comprehensive reviews Mondorf and Plank, (2024); Prasad et al., (2023); Golovneva et al., (2022) have further explored the reasoning behavior of LLMs beyond simple accuracy metrics, highlighting a prevalent concern: LLMs often rely on surface-level correlations rather than genuine logical inference. For instance, models might associate certain keywords or phrasing patterns with specific solution templates. CoCC directly tests this dependency by altering these patterns through methods like Storytelling, Gamification, and Domain Shift, all while maintaining the underlying semantic integrity of the task. This reliance on potentially superficial cues, even with advanced prompting techniques, underscores the need for evaluations like ours. In this work, we therefore aim to evaluate reasoning through more nuanced behavioral analyses rather than just final answer correctness, focusing on behavioral consistency under controlled linguistic and structural variations.
@@ 2.2 LLM for Code Generation
While these methods are valuable, metrics like pass@k primarily assess functional correctness on standard problem formulations, and LLM-based evaluators like CODEJUDGE, though assessing semantic correctness, may not fully capture how models respond when the presentation of the problem itself is adversarially yet semantically perturbed. Our study aligns with the goal of robust evaluation but extends it by focusing on the stability of the reasoning process leading to code generation when faced with such input variations. CoCC specifically assesses the impact of these adversarial modifications on the reasoning tasks inherent in code generation, probing whether the logical process itself is resilient.
@@ 2.3 Reasoning LLM Attacks
Adversarial attacks on code generation models have also been investigated. The DeceptPrompt framework Wu et al., (2023) reveals that benign-looking natural language modifications can induce LLMs to generate insecure or incorrect code, exposing vulnerabilities in real-world applications. While DeceptPrompt reveals vulnerabilities leading to insecure code, CoCC explores a complementary set of semantically-focused perturbations (e.g., Storytelling, Gamification, Domain Shift, Distracting Constraints). These are designed not primarily to induce insecurity, but to test the stability and fidelity of the underlying reasoning chain required for correct code generation when familiar problem cues are altered. Our study examines how these adversarial modifications—ranging from narrative restructuring to the injection of logically inert information—impact reasoning models. A key distinction in our methodology is the emphasis on perturbations that, while adversarial in their structure or framing, are designed to be semantically faithful to the original problem’s core logic. This allows us to isolate reasoning failures that are not due to misunderstanding a completely different task, but rather to an inability to reason robustly when the problem presentation changes.
In summary, prior work has established the potential of LLMs in complex reasoning and code generation, yet has also consistently highlighted their brittleness and potential over-reliance on learned superficial patterns. Existing evaluation methods and attack strategies provide valuable insights into model capabilities and vulnerabilities. However, a gap remains in systematically understanding how LLMs’ reasoning chains for code generation withstand diverse, semantically-grounded perturbations that alter problem presentation without changing the core logic. CoCC addresses this by introducing a novel framework and a suite of perturbation techniques specifically designed to probe these nuanced aspects of reasoning robustness in the domain of code generation.
@@ 3.2 Perturbation Methods
This transformation wraps the task in a challenge-based format (e.g., player or robots solving the task) while keeping the logic intact. Similar to storytelling, it introduces anthropomorphic or thematic context, drawing inspiration from instruction-tuning setups that present tasks in game-like scenarios (Wu et al.,, 2023). We evaluate whether such gamified inputs boosts engagement or distracts models from formal reasoning (see Figure 4).
@@
Inspired by the symbolic injection strategies in GSM-Symbolic (Mirzadeh et al.,, 2024), this method appends irrelevant but plausible-sounding constraints (e.g., ”input is a palindrome”) to the problem. While logically inert, these constraints test the model’s ability to filter out irrelevant linguistic noise. Unlike classic adversarial examples, these do not change task correctness but increase input ambiguity (see Figure 6).
@@
This strategy replaces tokens and terminology from one context to another (e.g., ”array of integers: becomes ”daily revenue”) or replacing terms with domain-specific equivalents (e.g., convert numbers to elevator floors, jobs, floors, days) (see Figure 7).
@@ 3.3 Logical Preservation Scoring
To quantify the degree of semantic deviation induced by each transformation, we developed a Logical Preservation Score. We employed the Claude-3.7-Sonnet model as an automated evaluator to ensure consistent and scalable scoring across all 700 perturbed instances. This approach mitigates potential human rater fatigue and subjectivity. The scoring is guided by a detailed rubric-based prompt (see Appendix F, Figure 10) that instructs the model to rate preservation on a scale of 1-10, where 10 indicates perfect logical alignment and 1 indicates a complete semantic shift. These scores are used in later analysis to understand how semantic drift correlates with performance degradation.
@@ 5.1 Clean Performance Overview
We begin by benchmarking all models on the original (unmodified) version of 100 LeetCode-style problems (Table 2). As expected, models optimized for instruction-following and reasoning, such as Gemini-2.5-Flash (95.0% overall) and Claude-3.7-Sonnet (90.0%), outperform code-specific models like Qwen2.5-Coder (59.0%) and DeepSeek-Coder-33B (54.0%). ON average, performance declines sharply from Easy (77.6%) to Hard (38.3%) tasks, reflecting the inherent challenge of generalization across reasoning complexity.
@@ 5.2 Perturbation Resilience and Gains
Table 3 reveals that semantic perturbations do not uniformly degrade performance. On the contrary, several perturbation types — particularly Storytelling and Gamification — consistently improve performance, especially for models tuned for human-aligned reasoning.
@@ 5.3 Semantic Robustness
In Figure 3, we plot average model accuracy against our logical preservation scores (Table 1. A positive, albeit nonlinear, correlation emerges: Storytelling (score: 8.89) and Gamification (8.01) lead to improved or stable performance. Distracting Constraints (3.82) and Negation Objective (1.83) consistently degrade it. However, we noticed that performance does not degrade proportionally with logical drift. Some logically preserved rewrites improve accuracy, while a single line of structurally trivial edits (like Distracting Constraints) dramatically hurt it. This suggests that cognitive alignment — not logical alignment alone — governs LLM success. LLMs are more sensitive to semantic distractors than to narrative abstraction.
@@ 5.5 Difficulty Analysis
As shown in Table 4, models often retain high accuracy on Easy problems across all attacks, but break under Hard tasks even for high-preservation rewrites. This confirms that reasoning complexity, not just input phrasing, lead to model breakdown. However, storytelling notably boosts hard-problem performance for Gemini-2.5 (+11.7%) and LLaMA3.1 (+23.5%), indicating that human-friendly reformulations may facilitate deeper reasoning under challenge.
@@ 5.6 Errors through Prompt Clarity
This highlights a dual nature of perturbations: when they disrupt expected patterns without adding structure, they degrade performance; but when they elaborate or scaffold problem context, they may improve reasoning as we see above for certain models in Table 5. Future models may benefit from tuning not only on correct answers, but on uncertainty aware reasoning paths.
@@ 5.7 Comment-Driven Reasoning Behaviors
A notable behavioral pattern emerged when analyzing Claude-3.7-Sonnet’s performance under the Gamification prompt variant. While Claude-3.7 demonstrates strong performance under clean conditions (90.0% overall accuracy; Table 2), its accuracy drops dramatically by –40.0% under Gamification (Table 3). Manual inspection of model outputs (Table  6) revealed a consistent shift in behavior: Claude’s code responses on unperturbed prompts routinely included inline comments—clarifying control flow logic, edge case handling, or the purpose of specific variables. In contrast, responses under gamified prompt formulations lacked such explanatory commentary.
@@ 6.1 Methods
This method fully inverts the core task (e.g., ”maximize” $\rightarrow$ ”minimize” or ”include” $\rightarrow$ ”exclude”). A part of the transformation prompt asks the LLM to: ”rewrite it with the opposite objective … ensure the examples and output are updated accordingly: (see Figure 8). While the input, output, explanation and constraints are preserved, the solution logic must change, requiring the model to recompute from a new problem intent. The preservation score was recorded as 1.83 based on Claude-3.7.
@@
This variant applied milder semantic reversals (e.g., asking for ”non-decreasing” instead of ”increasing”) without flipping the task entirely. The instruction explicitly states: ”Do NOT fundamentally change the problem’s task … adjust the examples minimally” (see Figure 9). This design tests whether models are robust to subtle linguistic tweaks that alter logic while retaining solvability. The preservation score was recorded as 2.56 based on Claude-3.7, a little higher than the hard negation.
@@ 6.2 Results
Table 7 summarizes the impact of negation-based perturbations. Hard negation proves highly disruptive, causing average accuracy drops of over 50% across all difficulty levels. For instance, Gemini-2.5-Flash drops from 95.0% to 42.6%, and Claude-3.7 falls from 90.0% to 15.0%. Even instruction-following models struggle to adapt, as reasoning reversal—not just language adaptation—is required. In contrast, soft negation shows more variability, with some models like LLaMA-3.1 improving from 19.0% to 29.4%. However, others like DeepSeek-14B and Claude-3.7 still suffer large drops, showing that even subtle logical shifts can disrupt reasoning.
@@ 8 Conclusion
In this work, we presented CoCC, a systematic framework for evaluating reasoning robustness in LLMs under adversarial prompting in code generation tasks. By introducing semantically plausible but logically challenging perturbations such as storytelling, gamification, distracting constraints, and negation objectives, we revealed key vulnerabilities in current state-of-the-art models. Through comprehensive experiments across nine LLMs and 700 perturbed instances, we found that while some perturbations surprisingly improved model performance, others, especially those involving subtle semantic drift, led to significant reasoning failures.